Which expression must be added to 3x - 7 to equal 0? (A) 0 (B) 3x + 7 (C) -3x + 7 (D) -3x - 7
Answer by Theo(13342)

3x - 7 = 0
3x = 7
x = 7/3
that just solves for x but doesn't answer the question.
your solutions are:
A.
adding 0 to 3x - 7 results in 3x - 7 so that isn't a solution.
B.
adding 3x + 7 to 3x + 7 results in 6x + 14 so that isn't a solution.
C.
adding -3x + 7 to 3x + 7 results in 14 so that isn't a solution.
D.
adding -3x - 7 to 3x + 7 results in 0 (bingo !!!!!)
your equation becomes (3x+7) + (-3x-7) = 3x + 7 - 3x - 7 which results in 0 after you combine all like terms.